OpenStudy (anonymous):

A person earns $187.42 before deductions for 3 days' work. What is the rate of pay per hour assuming 8 hours of work per day?

OpenStudy (♪chibiterasu):

Assume hours is h and days is d. We will be setting up an equation to solve for the rate (r). \[r = \frac {187.42}{hd}\] \[r = \frac {187.42}{8\times3}\] \[r = \frac {187.42}{24}\] \[r \approx 7.80197\] This is only an approximate value since the decimal was rounded. So you could say that the person earns about $7.80 per hour.